# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v3

@@ 4.2 PuM-based IMPACT Covert-Channel Attack
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [52] that enables bulk data movement operations (e.g., bulk data copy and initialization) within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. [^2] There could be near-memory RowClone variants that move data without using operational principles of DRAM as well. However, in this attack, we assume a PuM-based RowClone operation. RowClone’s Fast Parallel Mode works by issuing back-to-back two activation commands: the first activation copies the content of a source row (i.e., src) into the row buffer. The second command connects the destination row (i.e., dst) to the bitlines. Since the row buffer already stored the src’s data, by the time the dst is activated, the data in dst is overwritten by src’s data.
@@
To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
@@
Attack Overview. PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. The sender transmits a message by executing a RowClone operation with three operands: (i) a source address range, (ii) a destination address range, and (iii) a mask to select which parts of the ranges will be copied. Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks. The receiver repeatedly issues RowClone operations and measures the latency of these operations. This way, if the sender issues a RowClone operation to a DRAM bank, it induces contention, and the receiver detects a higher latency for its RowClone operation, indicating the transmitted bit is logic-1. Otherwise, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
@@
Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
Figure 5: PuM covert channel attack flow.
@@
Step 1. First, the receiver initializes all DRAM banks by issuing a RowClone operation (Line 16). Then, the sender and the receiver synchronize on a barrier (Lines 10 and 17) which ensures that they start the communication turn at the same time.
Step 2. The sender uses a single RowClone operation to transmit an N-bit message (where N is the number of banks) to the receiver in parallel and uses the mask of the RowClone operation to encode the message (Lines 11-12). The sender encodes the message by setting the corresponding bits in the mask to 1 or 0 ( 1, 2). To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4). The sender synchronizes with the receiver using a barrier operation (Lines 13 and 18).
@@ 1 Memory Controller:
3 for each DRAM bank do:
4 if mask[bank] == 1 then:
5 row_clone(bank, src, dst);
@@ 9 Sender:
10 barrier_1();
11 mask = init(message);
@@
13 barrier_2();
14
@@ 15 Receiver:
17 barrier_1();
18 barrier_2();
19 for each DRAM bank do:
@@
22 row_clone(src, dst, mask);
@@
25 if latency > threshold:
@@
Step 3. The receiver in IMPACT-PuM detects contention in DRAM banks by issuing RowClone operations and measuring their latency. The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24). If the sender induces contention in a DRAM bank, the receiver observes high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26). Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
Advantage over IMPACT-PnM. In IMPACT-PuM, the sender can use a single RowClone operation to transmit an N-bit message to the receiver in parallel, where N is the number of DRAM banks. Thus, IMPACT-PuM requires less computational resources to transmit the same amount of information compared to IMPACT-PnM. For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel. In contrast, in IMPACT-PnM, the sender needs to execute one PEI per bit.
4.3 PnM-based Side-Channel Attack on Read Mapping
@@
Read Mapping. DNA is one of the unique identifiers of an individual [73], such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity [70, 71]. Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information. Sequencing the DNA of a biological sample produces short DNA fragments called reads. The reads are then processed by read mapping, which is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
@@
Read mapping (RM) has two key steps: (i) seeding and (ii) alignment. During the seeding step, RM identifies the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome. After seeding, RM aligns the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching position among all possible locations in the alignment step.
@@
RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 96, 97, 98, 99, 100, 101, 102]. In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the sample genome.
@@
Attack Overview. We assume that the victim is running a RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [103]). The read mapping tool constructs a hash table that contains information about the seed locations in the reference genome, and allows each user in the system to probe the hash table via queries during read mapping. We assume the hash table is distributed across multiple DRAM banks. This assumption is realistic as many modern DRAM address mapping schemes (e.g., [104, 105, 106, 107]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [108, 109].
@@
Fig. 6 shows the high-level overview of the attack. The victim application extracts the seeds from the reads and offloads the corresponding seeding and alignment step to the PiM-enabled system  1. During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2. The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3. If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome. As shown in previous works on DNA imputation [110, 111, 112, 113], the attacker can use the leaked information in a completion attack [71] to infer properties about some regions of the private sample genome  4.
@@
Figure 6: High-level overview of IMPACT attack model on PiM-enabled RM.
@@ 5 Evaluation
5.1 Methodology
Covert Channel Threat Model. We assume a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to user space applications without any restrictions. The receiver has access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access latencies. For IMPACT-PnM, the sender and the receiver co-locate data in the same DRAM banks. For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity: when the sender issues a RowClone operation that spans across multiple banks, all individual bank-level RowClone operations need to be completed before another DRAM operation is executed.
Side Channel Threat Model. We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use an RM implementation based on minimap2 [103], and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system. We assume that the alignment step includes chaining. We use the human genome as the reference and compare it against synthetic sample genomes. We experiment with multiple seed sizes and report the best-performing selection.
Proof-of-Concept. We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [114]. [^3] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions to measure execution cycles from the user space.
Table 2 shows the configuration of our simulation setup. Before launching the attack, the sender and the receiver threads warm up by executing both attacks to avoid compulsory DRAM accesses for instructions or page table walks. For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of seven different covert-channel attacks.
(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [68] which uses clflush instructions to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [68] which uses cache eviction sets to send memory requests directly to main memory.
(iii) Streamline: State-of-the-art cache covert channel proposed in [115] which enables a flushless communication.
(iv) DMA Engine: A row buffer-based attack that uses the DMA engine to bypass the caches. We assume a powerful attacker can infer the DRAM access time directly from the user space by ignoring context switch and OS-related latencies.
(v) PnM-OffChip: A PnM-based attack in a PnM architecture with locality monitor and a state-of-the-art off-chip predictor [116]. If the data is predicted to have high locality, the computation is performed by the CPU instead of the near-memory compute units, leading to reduced attack throughput.
(vi) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
(vii) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack. To calculate their throughput, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack. We validate the model against the real system throughput of DRAMA and Streamline (e.g., Streamline [115] reports a maximum communication throughput of 1.8 Mb/s and we model the upper bound as 2.7 Mb/s for the smallest LLC).
Noise Sources in Simulation Setup and Error Rate. We simulate hardware prefetchers and page table walkers to induce noise in the simulated system. We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating the error rate in our measurements).
Table 2: Simulation Configurations
CPU 4-core, 4-way, OoO x86 core, 2.6GHz
MMU L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access
L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way assoc, 4-cycle access latency, LRU, IP-stride prefetcher [117]
L2 Cache 1 MB, 16-way assoc, 12-cycle latency, SRRIP [118], Streamer [119]
L3 Cache 2 MB/core, 16-way assoc, 32-cycle latency, SRRIP [118]
Main Memory DDR4-2400, 16 Banks, 4 Ranks, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
@@ 5.2 Validation of Proof-of-Concept
We first demonstrate the PoC results and how we successfully mount the attack in the simulated system. Fig. 7 shows the latency (in cycles) measured by the receiver after executing (a) a PEI instruction and (b) a RowClone operation, one for each of the 16 banks assuming the message is 16-bits large, in IMPACT-PnM and IMPACT-PuM, respectively. We observe that the receiver can successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM. The receiver can decode the complete message using a threshold value of 150 cycles as follows. If the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0, otherwise (i.e., row buffer conflict) it decodes the bit as logic-1.
@@
Figure 7: Latency (in cycles) measured by the receiver to execute (a) a PEI instruction and (b) a RowClone instruction for each bank in order to decode a 16-bit message.
5.3 Covert-Channel Attack Throughput Analysis

### v4

@@ 4.2 PuM-based IMPACT Covert-Channel Attack
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [75] which enables bulk data copy and initialization operations within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. RowClone’s Fast Parallel Mode works by issuing two activation commands back-to-back: the first activation copies the content of a source row (i.e., src) into the row buffer. The second connects the destination row (i.e., dst) to the bitlines. Since the row buffer stores the src’s data after the first activation, by the time the dst is activated, the data in dst is overwritten by src’s data.
@@
To execute RowClone, we assume that the userspace application specifies 1) a source virtual address range, 2) a destination virtual address range, and 3) a mask to copy specific parts of the source range to the destination range. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
@@
Attack Overview. Our PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. Before the attack, the sender and the receiver co-locate their data in the same set of DRAM banks. The sender transmits a message to the receiver in multiple batches of M bits. To do so, the sender issues RowClone operations by carefully selecting 1) source and destination address ranges corresponding to physical pages spanning the selected set of DRAM banks, and 2) a mask to cause interference only in selected DRAM banks. The sender and receiver processes synchronize with barriers similar to IMPACT-PnM (§4.1)
@@
Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and the steps for establishing the covert channel between the sender and the receiver.
Figure 5: PuM covert-channel attack flow.
@@
Step 1. First, the receiver initializes the DRAM banks by issuing a RowClone operation (Line 25, 1).
Step 2. The sender transmits the message by dividing the message into M-bit batches. Then, the sender maps each bit of a batch to a separate DRAM bank and creates a mask for the batch (Line 18), determining which banks will be executing a RowClone operation (Line 18). By using a mask, the sender uses only a single RowClone operation to transmit an M-bit batch of the message (where M is the number of banks) (Line 20, 2). The memory controller issues individual RowClone operations to each DRAM bank based on the specified mask (Lines 1-9). To send a 1, the sender triggers RowClone in the corresponding DRAM bank and activates a different row from the receiver’s row in the bank (Lines 4-7, 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not interfere with the receiver’s activated row (Lines 8-9, 4).
Step 3. The receiver detects interference in DRAM banks using RowClone operations as follows. The receiver issues one RowClone operation for each DRAM bank to precisely measure the time it takes to complete the operation in each bank (Lines 29-33). To detect if the sender activated another row, the receiver uses its initial destination address range as the source (i.e., swapping the direction of the copy operation) (Line 31). If the sender induces interference in a DRAM bank, the receiver observes high latency for the RowClone operation, indicating the corresponding bit in the message is set to 1 (Lines 34-35, 5). Otherwise, the receiver detects a fast RowClone operation, which indicates that the bit is set to 0 (Lines 36-37, 6).
@@ 1 Memory Controller:
3 for each DRAM bank:
4 if mask[bank] == 1:
5 // copy the corresponding src row
6 // to the destination row
7 row_clone(src, dst, bank);
@@ 11 Sender:
12 message[0:N-1]; // N-bit message
13 batch_size = M; // M-bit batch, where M<N
14 start = 0;
15 for every batch:
16 end = start + batch_size;
17 // create mask for the batch
18 mask = message[start:end];
19 // issue a RowClone operation using the mask
@@
21 start = end; // update start for the next batch
22 memory_fence(); // complete the batch
23
@@ 24 Receiver:
26 for every batch:
27 for each DRAM bank:
@@
30 // issue RowClone to one bank at a time
31 row_clone(dst, src, mask);
@@
34 if latency > THRESHOLD: // row buffer conflict
@@
Advantage over IMPACT-PnM (§4.1). In IMPACT-PuM, the sender can use a single RowClone operation to transmit an M-bit message to the receiver in parallel, where M is the number of DRAM banks. Thus, IMPACT-PuM requires fewer computational resources to transmit the same amount of information compared to IMPACT-PnM. For example, using a single thread, the sender can activate all M DRAM banks and transmit M bits of information in parallel. In contrast, in IMPACT-PnM (§4.1), the sender needs to execute one PEI for every bit.
4.3 PnM-based Side-Channel Attack on Genomic Read Mapping
@@
Genomic Read Mapping. DNA is a unique identifier of an individual [103], such that even a small fraction of DNA can include sufficient information for linking a DNA sample to the owner’s identity [100, 101]. Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information. Sequencing the DNA of a biological sample produces short DNA fragments called reads [141]. The reads are then processed by read mapping [34, 99, 36, 100, 101, 102, 103], which is a fundamental task performed to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
@@
Read mapping (RM) has two key steps: (i) seeding [142] and (ii) alignment [143, 144]. Fig. 6 shows a high-level overview of RM. During the seeding step, RM identifies the possible locations of the reads in the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) ( 1) and probing a hash table that is built from the reference genome ( 2). The hash table provides potential locations for the seed in the reference genome ( 3). In the alignment step, RM aligns the reads with their candidate regions in the reference (i.e., potential reference segments) using dynamic programming methods [99, 100] to find each read’s best matching position among all possible locations ( 4).
Figure 6: DNA Read Mapping Operation: Seeding and Sequence Alignment.
@@
RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [37, 145, 146, 40, 147, 148, 149, 150, 35, 151, 152]. In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the query genome.
@@
Attack Overview. We assume that the victim is running an RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [153]). The read mapping tool constructs a hash table that contains information about the seed locations in the reference genome, and allows each user in the system to probe the hash table via queries during read mapping. We assume the hash table is distributed across multiple DRAM banks. This assumption is realistic, as many modern DRAM address mapping schemes (e.g., [154, 155, 156, 157]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [158, 159].
@@
Fig. 7 shows the high-level overview of the attack. The hash table is distributed across banks as explained above ( 1). During the attack, the victim application extracts seeds from the reads and offloads the corresponding seeding and alignment steps to the PiM-enabled system ( 2). During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-M of the hash table) ( 3). The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table ( 4). If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome. As shown in previous works on DNA imputation [160, 161, 162, 163], the attacker can use the leaked information in a completion attack [101] to infer properties about some regions of the private query genome ( 5).
@@
Figure 7: High-level overview of the IMPACT attack model on PiM-enabled genomic read mapping.
5 Methodology
5.1 Threat Models
Covert Channel Threat Model. We assume a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to userspace applications without any restrictions. The receiver has access to cpuid and rdtscp instructions [121], enabling high-precision measurement of memory access latencies. For IMPACT-PnM, the sender and the receiver co-locate data in the same set of DRAM banks. For IMPACT-PuM, the sender allocates two virtual address ranges whose corresponding physical pages span across the same set of DRAM banks.
Side Channel Threat Model. We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use a genomic read mapping implementation based on minimap2 [153], and we modify the implementation to offload the seeding and alignment steps to the PiM-enabled system [67]. We assume that the alignment step includes chaining [99, 153]. We use the human genome as the reference and compare it against synthetic query genomes. We experiment with multiple seed sizes and report the best-performing selection.
5.2 Evaluation Methodology
5.2.1 Proof-of-Concept
We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [164]. [^2] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions [121] to measure execution time.
Table 2 shows the configuration of our simulation setup. Before launching the attack, the sender and the receiver threads warm up to avoid compulsory DRAM accesses for instructions or page table walks. For the PnM system [67], we model the additional latency of a PEI (e.g., latency of accessing additional system structures) as $3$ cycles [67].
Table 2: Simulated System Configuration
CPU 4-core, 4-way issue, OoO x86 core, 2.6GHz
MMU L1 DTLB (4KB): 64-entry, 4-way, 1-cycle, L1 DTLB (2MB): 32-entry, 4-way, 1-cycle, L2 TLB: 1536-entry, 12-way, 12-cycle
L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way, 4-cycle, LRU, IP-stride prefetcher [165]
L2 Cache 2 MB, 16-way, 16-cycle, SRRIP [166], Streamer prefetcher [167]
L3 Cache 2 MB/core, 16-way, 50-cycle, SRRIP [166]
Main Memory DDR4-2400, 16 Banks, 4 bank groups, 1 Rank, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
5.2.2 Comparison Points for Covert-Channel Attacks
To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of five different covert-channel attacks.
(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [98] which uses clflush instructions [121] to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [98], which uses cache eviction sets [122] to send memory requests directly to main memory.
(iii) DMA Engine: A row buffer-based attack that uses the DMA engine [123, 124] to bypass the cache hierarchy.
(iv) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
(v) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
5.2.3 Noise Sources in Simulation Setup and Error Rate
We simulate hardware prefetchers [167, 165] and page table walkers [121] to induce noise in the simulated system. We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating error rate in our measurements).
@@ 6.1 Validation of Proof-of-Concept
We first demonstrate proof-of-concept attacks for IMPACT-PnM and IMPACT-PuM. We send 16-bit messages with a batch size of 16 (i.e., using 16 banks in parallel). Fig. 8 shows the latency (in cycles) measured by the receiver during transmission of example messages. The y-axis shows the measured latency for each (a) PEI and (b) RowClone operation for IMPACT-PnM and IMPACT-PuM, respectively. The x-axis shows the transmitted bit values where each bit is transmitted using a different bank. We make two key observations. First, for both attacks, the receiver successfully determines the transmitted bit values by detecting the row buffer conflicts caused by the sender. Second, the receiver decodes the complete message using a threshold value of 150 cycles. If the measured latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0; otherwise (i.e., row buffer conflict), it decodes the bit as logic-1.
@@
Figure 8: Latency (in cycles) measured by the receiver to execute (a) a PEI and (b) a RowClone operation for each bank to decode a 16-bit message.
6.2 Covert Channel Throughput Analysis
Fig. 9 shows the throughput achieved by IMPACT-PnM, IMPACT-PuM and state-of-the-art main memory-based covert-channel attacks in modern systems with increasing LLC sizes. The x-axis shows the increasing LLC size (from 1 MB to 128 MB), and the y-axis shows the leakage throughput of the covert channels. For all attacks, one sender and one receiver process run concurrently and transmit a fixed-length message.